# UltraWiki: Ultra-fine-grained Entity Set Expansion with Negative Seed Entities

### v1

@@ 6.1. Experiment Setup
where $Q$ stands for the set of all queries, $\mbox{AP}_{K}\left(L_{q},\mathcal{P}\right)$ denotes the average precision at position $K$ with the ranked list $L_{q}$ and ground-truth list $\mathcal{P}$. P@$K$ is the precision of the top-$K$ entities.
@@
Table 2. Main experiment results. We report the MAP and P scores corresponding to Pos/Neg/Comb. The highest average score is bolded and the next highest average score is underlined.
Method Type Method Metric Type MAP P Avg
@@
Probability Based SetExpan Pos $\uparrow$ 13.41 11.83 10.66 10.79 20.10 19.97 21.88 26.39 16.88
Neg $\downarrow$ 4.06 3.77 3.71 4.20 7.66 8.10 10.92 17.44 7.48
Comb $\uparrow$ 54.67 54.03 53.48 53.30 56.22 55.93 55.48 54.48 54.70
CaSE Pos $\uparrow$ 16.72 13.74 11.60 10.91 24.58 23.28 26.86 30.53 19.78
Neg $\downarrow$ 5.33 4.32 3.63 3.50 10.22 10.10 12.79 15.96 8.23
Comb $\uparrow$ 55.69 54.71 53.99 53.70 57.18 56.59 57.03 57.28 55.77
Retrieval Based CGExpan Pos $\uparrow$ 21.64 19.72 19.11 20.22 30.61 31.24 38.39 50.03 28.87
Neg $\downarrow$ 6.15 6.54 8.03 9.96 12.29 16.37 27.38 41.72 16.06
Comb $\uparrow$ 57.75 56.59 55.54 55.13 59.16 57.44 55.50 54.15 56.41
ProbExpan Pos $\uparrow$ 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
Neg $\downarrow$ 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
Comb $\uparrow$ 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
Generation Based GPT4 Pos $\uparrow$ 37.20 35.37 35.49 35.59 47.12 48.87 55.31 62.22 44.65
Neg $\downarrow$ 6.04 6.61 8.03 8.35 10.40 15.06 24.57 33.63 14.09
Comb $\uparrow$ 65.58 64.38 63.73 63.62 68.36 66.90 65.37 64.29 65.28
Retrieval Based (Ours) RetExpan Pos $\uparrow$ 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
Neg $\downarrow$ 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
Comb $\uparrow$ 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
+ Contrast Pos $\uparrow$ 47.45 44.68 43.63 44.20 59.83 62.02 69.36 77.92 56.14
Neg $\downarrow$ 8.02 8.98 10.89 13.05 14.83 21.23 35.47 55.12 20.95
Comb $\uparrow$ 69.72 67.85 66.37 65.57 72.50 70.39 66.95 61.40 67.59
+RA Pos $\uparrow$ 44.74 42.23 40.66 40.27 58.31 61.36 66.58 73.97 53.52
Neg $\downarrow$ 5.53 5.88 6.56 7.44 11.70 16.70 25.81 37.05 14.58
Comb $\uparrow$ 69.60 68.18 67.05 66.42 73.31 72.33 70.39 68.46 69.47
Generation Based (Ours) GenExpan Pos $\uparrow$ 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
Neg $\downarrow$ 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
Comb $\uparrow$ 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
+ CoT Pos $\uparrow$ 50.39 47.80 43.67 40.06 62.74 64.45 64.06 60.38 54.19
Neg $\downarrow$ 7.79 9.29 8.15 6.89 15.97 22.66 23.52 21.90 14.52
Comb $\uparrow$ 71.30 69.25 67.76 66.58 73.39 70.90 70.27 69.24 69.84
+ RA Pos $\uparrow$ 48.68 46.26 43.80 41.27 61.12 62.64 66.49 65.65 54.49
Neg $\downarrow$ 7.16 8.54 8.63 7.70 14.79 21.18 26.67 26.84 15.19
Comb $\uparrow$ 70.76 68.86 67.58 66.78 73.16 70.73 69.91 69.41 69.65
6.2. Main Experiments
The results of the main experiment are presented in Table 2, from which we can observe that:
@@ 6.2. Main Experiments
The ablation experiments presented in Table 3 further validate our hypothesis regarding the factors why RetExpan and GenExpan surpass the previous state-of-the-art models. Both methods exhibit remarkable degradation after removing the corresponding modules.
(3) The simple retrieval augmentation strategy yielded remarkable results for both RetExpan and GenExpan. Particularly for RetExpan, the strategy improves the comprehensive Comb metrics by 4.11 points on average. The strategy primarily optimizes the Neg class metrics, leading to a significant reduction in the intrusion of negative target entities. However, its effect on Pos class metrics exhibits instability. For instance, the PosMAP@50 and PosMAP@100 of RetExpan+RA are instead lower compared to the original RetExpan. This phenomenon can be attributed to the fact that the extra introductions corresponding to each entity are currently fixed and do not adapt to changes in the entity context. Consequently, this increases the homogeneity of the entity embedding between irrelevant entities (neither positive nor negative target entities) and positive seed entities under the same fine-grained semantic class to some extent. Our point is further supported by the observation that RetExpan’s MAP@100 for fine-grained semantic classes grows from 82.08 to 87.17 after the introduction of retrieval augmentation strategy. This suggests that a dynamic and more fine-grained knowledge retrieval strategy needs further investigation.
(4) Among all the enhancement strategies, ultra-fine-grained contrastive learning provides the most substantial improvement in the Pos class metrics, averaging 4.10 points (from 52.04 to 56.14). Conversely, contrastive learning brings smaller gains in Neg class metrics. This discrepancy is primarily attributed to the implementation of contrastive learning, where both positive and negative target entities are pulled away from normal negative entities originating from other semantic classes to ensure underlying semantics. This dilutes the pulling away of positive and negative target entities from each other, as they are both in the denominator of the contrastive loss and are assigned the same weight.
@@
(5) The conventional probability-based methods SetExpan and CaSE attain quite low scores on both NegMAP and NegP metrics, which indicates that negative target entities are not excessively introduced. However, this does not imply that these methods exhibit a high degree of negative semantic awareness. The Pos class metrics of SetExpan and CaSE are similarly low, indicating their limited understanding of fine-grained semantic classes overall. Consequently, they struggle to recall entities with the given seed entities, leading to low scores in both Pos and Neg class metrics. We further evaluated the MAP@100 of CaSE at the fine-grained semantic class level, which only reaches 21.43, significantly lower than other methods such as 82.08 of RetExpan.
(6) Without further fine-tuning and modifications to the model structure, GPT-4 achieves excellent results, which proves that generative LLMs have great potential for addressing Ultra-ESE tasks. However, GPT-4 still fails to outperform the GenExpan, which is also generative and based on smaller LLaMA-7B. Analysis of cases reveals that there are two primary problems with GPT-4. On the one hand, it performs poorly on long-tail problems, such as U.S. national monuments, mobile phone brands, which contain a considerable number of low-frequency entities with limited information available on the Internet.GPT-4 only achieves single-digit PosMAP on these semantic classes. In contrast, GenExpan performs better, benefiting from the given contextual corpus and targeted retrieval augmentation strategies. On the other hand, GPT-4 is prone to haphazardly generate non-existent entities (e.g., fake mobile phone brands), which is referred to hallucination problem in recent work(Zhang et al., 2023a; Ji et al., 2023). We are currently unable to solve this issue by simple output post-processing. While the results of GPT-4 are impressive, there remains a lot of space for improvement.
Table 3. Ablation experiments for each module of RetExpan and GenExpan. Each module contributes to the overall performance.
@@
6.3. Analysis of Negative Entities
@@
Whether the introduction of negative seed entities is effective for representing ultra-fine-grained semantic classes? The negative seed entity-based entity re-ranking modules were removed from RetExpan and GenExpan. Thanks to the high scalability, it was also integrated into ProbExpan. Results of the three comparison experiments are presented in Table 5. We can observe that:
@@
(2) Further analyzing the Pos and Neg metrics, we observe that the addition of the entity re-ranking module to ProbExpan conversely resulted in a notable decline in the PosP metrics. It’s essential to clarify the distinction between the MAP and P metrics: P@K is concerned solely with the number of target entities in the top-K entity list, regardless of their positions, while MAP@K is rank-aware, with higher values indicating that target entities are positioned closer to the top of the list. Consequently, this indicates that the incorporation of negative seed entities in ProbExpan makes a mixed impact, propelling some positive target entities to higher ranks and simultaneously causing others to be ranked lower.
(3) For RetExpan, negative seed entities have a similar impact on metrics with different K values, whereas for GenExpan, negative seed entities have a greater impact on metrics with a smaller K. This may be due to the fact that RetExpan is a one-time expansion (Kushilevitz et al., 2020; Mamou et al., 2018; Yu et al., 2019) model and GenExpan is an iterative expansion (Huang et al., 2020; Rong et al., 2016; Shen et al., 2017) model, so entities expanded later (corresponding to a larger K) are more likely to deviate from the original ground-truth semantic class. During these stages, re-ranking using negative seed entities also doesn’t work. Addressing the issue of semantic drift (Curran et al., 2007; McIntosh, 2010; Shi et al., 2014) in iterative expansion models like GenExpan is a longstanding challenge in ESE field.
Table 4. Comparison experiments when positive and negative attributes are the same and different.
Method Metric Type MAP Avg
@10 @20 @50 @100
$\mathbf{\mathcal{A}^{pos}=\mathcal{A}^{neg}}$
RetExpan Pos 43.14 41.74 41.93 42.89 42.43
Neg 5.54 6.44 8.71 11.56 8.06
Comb 68.80 67.65 66.61 65.67 67.18
RetExpan + Contrast Pos 49.68 48.73 48.77 49.50 49.17
Neg 5.76 6.51 8.64 11.62 8.13
Comb 71.96 71.11 70.06 68.94 70.52
RetExpan + RA Pos 46.29 44.03 43.46 43.61 44.35
Neg 5.45 5.86 6.84 7.56 6.43
Comb 70.42 69.08 68.31 68.03 68.96
$\mathbf{\mathcal{A}^{pos}\neq\mathcal{A}^{neg}}$
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.44
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan + Contrast Pos 46.44 43.69 42.51 43.14 43.95
Neg 8.22 9.25 11.17 13.38 10.51
Comb 69.11 67.22 65.67 64.88 66.72
RetExpan + RA Pos 44.74 42.23 40.66 40.27 41.98
Neg 5.53 5.88 6.56 7.44 6.35
Comb 69.60 68.17 67.05 66.42 67.81
Table 5. Ablation experiments on entity re-ranking module with negative seed entities.
@@
What are the roles played by negative seed entities? We first assessed the efficacy of RetExpan and its enhancement strategies by comparing scenarios where positive and negative attributes are identical or different. As shown in Table 4, we can find that:
(1) When positive and negative attributes are identical, there is no overlap between positive and negative target entities, which simplifies semantic classes and results in relatively high performance. In such cases, negative seed entities play a supportive role, primarily emphasizing user-concerned attributes and mitigating ambiguity. Conversely, when positive and negative attributes are different, the potential overlap between entities satisfying these attributes poses a significant challenge to semantic understanding, leading to lower overall performance. Here, negative seed entities play a crucial role in conveying unwanted semantics and are indispensable.
(3) When positive and negative attributes differ, the primary focus lies in evaluating the model’s capability to comprehend the semantics expressed by negative seed entities. We observe that retrieval augmentation demonstrates more effective in this situation. This phenomenon suggests that the understanding of negative seed entities demands more additional knowledge compared to positive seed classes. This notion is further verified in the analysis of chain-of-thought reasoning in Section 6.4.
What are the differences in challenges posed by semantic classes with varying numbers of positive and negative attributes? We analyze the performance of RetExpan on semantic classes characterized by varying quantities of positive and negative attributes in Table 6. To maintain control over variables, we exclusively present semantic classes for three combinations of attribute quantity: (1,1), (1,2), and (2,1). Imposing stricter constraints on the attributes results in a reduction in the number of associated target entities, rendering it more challenging to discern these entities accurately. Consequently, a decrease in the Pos (Neg) metrics is observed, correlating with an increase in the number of positive (negative) attributes. It’s important to emphasize that although the Neg metrics decrease with more negative attribute constraints, this does not imply a reduced difficulty in distinguishing negative target entities. Rather, identifying these negative target entities requires the model to possess more comprehensive attribute information about the negative seed entities.
Table 6. Comparison experiments on semantic classes with different numbers of positive and negative attributes.
@@
6.4. Analysis of Enhancement Strategy
Table 7. Ablation experiments on contrastive learning.
@@ 6.4. Analysis of Enhancement Strategy
6.4.1. Analysis of Contrastive Learning
We conducted ablation experiments on each part of the training data for ultra-fine-grained contrastive learning to explore their impact on overall performance. As illustrated in Table 7, based on the full training data, the last three rows sequentially remove the hard negative samples comprising pairs of positive and negative target entities, the normal negative samples consisting of pairs of positive and negative target entities along with entities from other semantic classes, and the positive samples composed of entity pairs within the same ultra-fine-grained semantic class.
It can be seen that hard negative samples consisting of pairs of positive and negative target entities contribute the most to the overall performance, which is in line with our intuition of proposing ultra-fine-grained contrastive learning. We set both —$L_{pos}$— and —$L_{neg}$— to 10 to mitigate the introduction of excessive noise. Consequently, it’s evident that this part of training data wields more influence when K is small. Exploring how to mitigate the noise introduced by automatic GPT-4 labeling when —$L_{pos}$— and —$L_{neg}$— increase is a promising direction for future exploration.

### v3

@@ VI-A Experiment Setup
where $Q$ stands for the set of all queries, $\mbox{PosAP}_{K}\left(L_{q},\mathcal{P}\right)$ denotes the average precision at position $K$ with the ranked list $L_{q}$ and ground-truth list $\mathcal{P}$. PosP@$K$ is the precision of the top-$K$ entities.
@@
TABLE II: We report the MAP and P scores corresponding to Pos/Neg/Comb on UltraWiki. The highest score is bolded.
Metric Type Method Type Method MAP P Avg
@@
Pos $\uparrow$ Probability Based SetExpan 13.41 11.83 10.66 10.79 20.10 19.97 21.88 26.39 16.88
CaSE 16.72 13.74 11.60 10.91 24.58 23.28 26.86 30.53 19.78
Retrieval Based CGExpan 21.64 19.72 19.11 20.22 30.61 31.24 38.39 50.03 28.87
ProbExpan 21.86 22.11 22.80 23.89 38.08 39.41 47.02 62.71 34.74
Generation Based GPT4 37.20 35.37 35.49 35.59 47.12 48.87 55.31 62.22 44.65
Retrieval Based (Ours) RetExpan 41.73 39.53 38.55 39.91 54.58 58.03 66.76 77.23 52.04
RetExpan + Contrast 47.45 44.68 43.63 44.20 59.83 62.02 69.36 77.92 56.14
Generation Based (Ours) GenExpan 46.79 45.00 42.89 40.80 59.77 62.15 66.26 66.57 53.78
GenExpan + CoT 50.39 47.80 43.67 40.06 62.74 64.45 64.06 60.38 54.19
Neg $\downarrow$ Probability Based SetExpan 4.06 3.77 3.71 4.20 7.66 8.10 10.92 17.44 7.48
CaSE 5.33 4.32 3.63 3.50 10.22 10.10 12.79 15.96 8.23
Retrieval Based CGExpan 6.15 6.54 8.03 9.96 12.29 16.37 27.38 41.72 16.06
ProbExpan 6.72 8.16 10.85 13.47 15.12 19.92 34.51 56.48 20.65
Generation Based
GPT4 6.04 6.61 8.03 8.35 10.40 15.06 24.57 33.63 14.09
Retrieval Based (Ours) RetExpan 8.77 9.04 10.65 13.29 16.44 21.04 34.78 56.54 21.32
RetExpan + Contrast 8.02 8.98 10.89 13.05 14.83 21.23 35.47 55.12 20.95
Generation Based (Ours) GenExpan 7.25 8.28 8.72 8.01 15.21 21.31 27.33 28.58 15.59
GenExpan + CoT 7.79 9.29 8.15 6.89 15.97 22.66 23.52 21.90 14.52
Comb $\uparrow$ Probability Based SetExpan 54.67 54.03 53.48 53.30 56.22 55.93 55.48 54.48 54.70
CaSE 55.69 54.71 53.99 53.70 57.18 56.59 57.03 57.28 55.77
Retrieval Based CGExpan 57.75 56.59 55.54 55.13 59.16 57.44 55.50 54.15 56.41
ProbExpan 57.57 56.98 55.97 55.21 61.48 59.75 56.25 53.12 57.04
Generation Based
GPT4 65.58 64.38 63.73 63.62 68.36 66.90 65.37 64.29 65.28
Retrieval Based (Ours) RetExpan 66.48 65.25 63.95 63.31 69.07 68.50 65.99 60.34 65.36
RetExpan + Contrast 69.72 67.85 66.37 65.57 72.50 70.39 66.95 61.40 67.59
Generation Based (Ours) GenExpan 69.77 68.36 67.08 66.40 72.28 70.42 69.47 68.99 69.10
GenExpan + CoT 71.30 69.25 67.76 66.58 73.39 70.90 70.27 69.24 69.84
VI-B Main Experiments
The results of the main experiment are presented in Table II, from which we can observe that:
@@ VI-B Main Experiments
The ablation experiments presented in Table III further validate our hypothesis regarding the factors why RetExpan and GenExpan surpass the previous state-of-the-art models. Both methods exhibit remarkable degradation after removing the corresponding modules.
(3) Comparing the two enhancement strategies, ultra-fine-grained contrastive learning provides the most substantial improvement in the Pos class metrics, averaging 4.10 points (from 52.04 to 56.14). Conversely, contrastive learning brings smaller gains in Neg class metrics. This discrepancy is primarily attributed to the implementation of contrastive learning, where both positive and negative target entities are pulled away from normal negative entities originating from other semantic classes to ensure underlying semantics. This dilutes the pulling away of positive and negative target entities from each other, as they are both in the denominator of the contrastive loss and are assigned the same weight.
@@
(4) The conventional probability-based methods SetExpan and CaSE attain quite low scores on both NegMAP and NegP metrics, which indicates that negative target entities are not excessively introduced. However, this does not imply that these methods exhibit a high degree of negative semantic awareness. The Pos class metrics of SetExpan and CaSE are low in similarity, indicating their limited understanding of fine-grained semantic classes overall. Consequently, they struggle to recall entities with the given seed entities, leading to low scores in both Pos and Neg class metrics. We further evaluated the MAP@100 of CaSE at the fine-grained semantic class level, which only reaches 21.43, significantly lower than other methods such as 82.08 of RetExpan.
(5) Without further fine-tuning and modifications to the model structure, GPT-4 achieves excellent results, which proves that generative LLMs have great potential for addressing Ultra-ESE tasks. However, GPT-4 still fails to outperform the GenExpan, which is also generative but based on smaller LLaMA-7B. Analysis of cases reveals that there are two primary problems with GPT-4. On the one hand, it performs poorly on long-tail problems, such as U.S. national monuments and mobile phone brands, which contain a considerable number of low-frequency entities with limited information available on the Internet. GPT-4 only achieves single-digit PosMAP on these semantic classes. In contrast, GenExpan performs better, benefiting from the given contextual corpus. On the other hand, GPT-4 is prone to haphazardly generate non-existent entities (e.g., fake mobile phone brands), which is referred to hallucination problem in recent work[37, 38]. We are currently unable to solve this issue by simple output post-processing. While the results of GPT-4 are impressive, there remains a lot of space for improvement.
TABLE III: Ablation experiments for each module of RetExpan and GenExpan. Each module contributes to the overall performance.
@@
VI-C Analysis of Negative Entities
@@
Whether the introduction of negative seed entities is effective for representing ultra-fine-grained semantic classes? The negative seed entity-based entity re-ranking modules were removed from RetExpan and GenExpan. Thanks to the high scalability, it was also integrated into ProbExpan. Results of the three comparison experiments are presented in Table IV. We can observe that:
@@
(2) Further analyzing the Pos and Neg metrics, we observe that the addition of entity re-ranking module to ProbExpan conversely resulted in a notable decline in PosP metrics. It’s essential to clarify the distinction between MAP and P metrics: P@K is concerned solely with number of target entities in the top-K entity list, regardless of their positions, while MAP@K is rank-aware, with higher values indicating that target entities are positioned closer to the top of the list. Consequently, this indicates that the incorporation of negative seed entities in ProbExpan makes a mixed impact, propelling some positive target entities to higher ranks and simultaneously causing others to be ranked lower.
(3) For RetExpan, negative seed entities have a similar impact on metrics with different K values, whereas for GenExpan, negative seed entities have a greater impact on metrics with a smaller K. This may be due to the fact that RetExpan is a one-time expansion [39, 40, 36] model and GenExpan is an iterative expansion [41, 42, 18] model, so entities expanded later (corresponding to a larger K) are more likely to deviate from the original ground-truth semantic class. During these stages, re-ranking using negative seed entities also doesn’t work. Addressing the issue of semantic drift [25, 43, 27] in iterative expansion models like GenExpan is a longstanding challenge in the ESE field.
TABLE IV: Ablation experiments on entity re-ranking module with negative seed entities.
@@
TABLE V: Comparison experiments when positive and negative attributes are the same and different.
Method Metric Type MAP Avg
@10 @20 @50 @100
$\mathbf{\mathcal{A}^{pos}=\mathcal{A}^{neg}}$
RetExpan Pos 43.14 41.74 41.93 42.89 42.43
Neg 5.54 6.44 8.71 11.56 8.06
Comb 68.80 67.65 66.61 65.67 67.18
RetExpan + Contrast Pos 49.68 48.73 48.77 49.50 49.17
Neg 5.76 6.51 8.64 11.62 8.13
Comb 71.96 71.11 70.06 68.94 70.52
$\mathbf{\mathcal{A}^{pos}\neq\mathcal{A}^{neg}}$
RetExpan Pos 41.73 39.53 38.55 39.91 39.93
Neg 8.77 9.04 10.65 13.29 10.44
Comb 66.48 65.25 63.95 63.31 64.75
RetExpan + Contrast Pos 46.44 43.69 42.51 43.14 43.95
Neg 8.22 9.25 11.17 13.38 10.51
Comb 69.11 67.22 65.67 64.88 66.72
TABLE VI: Comparison experiments on semantic classes with different numbers of positive and negative attributes.
@@
What are the roles played by negative seed entities? We first assessed the efficacy of RetExpan and its enhancement strategies by comparing scenarios where positive and negative attributes are identical or different. As shown in Table V, we can find that:
(1) When positive and negative attributes are identical, there is no overlap between positive and negative target entities, which simplifies semantic classes and yielding higher performance. In such cases, negative seed entities play a supportive role, primarily emphasizing user-concerned attributes and mitigating ambiguity. Conversely, when positive and negative attributes are different, the potential overlap between entities satisfying these attributes poses a significant challenge to semantic understanding, leading to lower overall performance. Here, negative seed entities play a crucial role in conveying unwanted semantics and are indispensable.
(3) When positive and negative attributes differ, the primary focus lies in evaluating the model’s capability to comprehend the semantics expressed by negative seed entities. We observe that contrastive learning demonstrates less effective in this situation. This phenomenon suggests that the understanding of negative seed entities demands more additional knowledge compared to positive seed classes. This notion is further verified in the analysis of chain-of-thought reasoning in Section VI-D.
What are the differences in challenges posed by semantic classes with varying numbers of positive and negative attributes? We analyze the performance of RetExpan on semantic classes characterized by varying quantities of positive and negative attributes in Table VI. To maintain control over variables, we exclusively present semantic classes for three combinations of attribute quantity: (1,1), (1,2), and (2,1). Imposing stricter constraints on the attributes results in a reduction in the number of associated target entities, rendering it more challenging to discern these entities accurately. Consequently, a decrease in the Pos (Neg) metrics is observed, correlating with an increase in the number of positive (negative) attributes. It’s important to emphasize that although the Neg metrics decrease with more negative attribute constraints, this does not imply a reduced difficulty in distinguishing negative target entities. Rather, identifying these negative target entities requires more comprehensive attribute information about negative seed entities.
VI-D Analysis of Enhancement Strategy
TABLE VII: Ablation experiments on contrastive learning.
@@ VI-D Analysis of Enhancement Strategy
VI-D1 Analysis of Contrastive Learning
We conducted ablation experiments on each part of the training data for ultra-fine-grained contrastive learning to explore their impact on overall performance. As illustrated in Table VII, based on the full training data, the last three rows sequentially remove the hard negative samples comprising pairs of positive and negative target entities, the normal negative samples consisting of pairs of positive and negative target entities along with entities from other semantic classes, and the positive samples composed of entity pairs within the same ultra-fine-grained semantic class.
It can be seen that hard negative samples consisting of pairs of positive and negative target entities contribute the most to the overall performance, which is in line with our intuition of proposing ultra-fine-grained contrastive learning. We set both $|L_{pos}|$ and $|L_{neg}|$ to 10 to mitigate the introduction of excessive noise. Consequently, it’s evident that this part of training data wields more influence when K is small. Exploring how to mitigate the noise introduced by automatic GPT-4 labeling when $|L_{pos}|$ and $|L_{neg}|$ increase is a promising direction for future exploration.